# On Scaling Up 3D Gaussian Splatting Training

### v1

@@ 4.1 Empirical Evidence of Independent Gradients
To see if the Independent Gradients Hypothesis holds in practice, we analyze the average per-parameter variance of the gradients in real-world settings. We plot the sparsity and variance of the gradients of the diffuse color parameters starting at pre-trained checkpoints on the “Rubble” dataset [29] against the batch size in Figure 5. We find that the inverse of the variance increases roughly linearly, then transitions into a plateau. We find this behavior in all three checkpoint iterations, representing early, middle, and late training stages. The initial linear increase of the precision suggests that gradients are roughly uncorrelated at batch sizes used in this work (up to 32) and supports the independent gradients hypothesis. However, it is worth noting that even though a single image has sparse gradients, when there are many images in a batch, the gradients overlap and become less sparse. They also become more correlated because we expect images with similar poses to offer similar gradients.
Figure 5: Gradients are roughly uncorrelated in practice. On the “Rubble” dataset [29], the inverse of the average parameter variance increases linearly, then rises to a plateau, suggesting that the gradients are roughly uncorrelated initially but become less so as the batch size becomes large. Averaged over 32 random trials.
(a) Learning rate scaling rules vs. BS invariance.
(b) Momentum scaling rules vs. BS invariance.
Figure 6: We plot the training trajectories of the diffuse color parameters on “Rubble”, when training with batch size $\in[4,16,32]$ using different learning rate and momentum scaling strategies. Cumulative weight updates using the square-root learning rate scaling rule (a, red curves) and exponential momentum scaling rule (b, red curves) maintain high cosine similarity to batch-size 1 updates and have norms that are roughly invariant to the batch size.
@@ 4.2 Empirical Testing of Proposed Scaling Rules
To empirically test whether the proposed learning rate and momentum scaling rules work well, we train the “Rubble” scene to iteration 15,000 with a batch size of 1. Then, we reset the Adam optimizer states and continue training with different batch sizes. We compare how well different learning-rate and momentum scaling rules maintain a similar training trajectory when switching to larger batch sizes in Figure 6. Since different parameter groups of 3D GS have vastly different magnitudes, we focus on one specific group, namely the diffuse color, to make the comparisons meaningful. Figure 6(a) compares three different learning rate scaling rules $\in$ [constant, sqrt, linear] where only our proposed “sqrt” holds a high update cosine similarity and a similar update magnitude across different training batch sizes. Similarly, 6(b) shows our proposed exponential momentum scaling rule keeps update cosine similarity higher than the alternative which leaves the momentum coefficients unchanged.
@@ 5 Evaluation
Our evaluation aims to demonstrate Grendel’s scalability, showing both that it can render high-resolution images from large scenes, and that its performance scales with additional hardware resources. And we include results from abalation study to demonstrate the importance of Grendel’s dynamic load balancing. The ablation study for the learning rate scaling strategies have already been discussed in 4.2, along with our analysis.
@@ 5.1 Setting and Datasets
Experimental Setup.
We conducted our evaluation in the Perlmutter GPU cluster [24]. Each node we used was equipped with 4 A100 GPUs with 40GB of GPU memory, and interconnected with each other using 25GB/s NVLink per direction. Servers were connected to each other using a 200Gbps Slingshot network.
Datasets.
We evaluate Grendel using the datasets and corresponding resolution settings shown in Table 1. Of these, Rubble and MatrixCity Block_All represent the large scale datasets that are out of reach for most existing 3DGS systems, while other datasets are commonly used in 3DGS papers. These datasets vary in area size and resolution to comprehensively test our system.
@@
Tanks & Temple [11] 2 $\sim$ 1K 251 to 301 1/8 of all images
DeepBlending [6] 2 $\sim$ 1K 225 to 263 1/8 of all images
Mip-NeRF 360 [1] 9 1080P 100 to 330 1/8 of all images
Rubble [29] 1 $4591\times 3436$ 1657 official test set
MatrixCity Block_All [15] 1 1080P 5620 official test set
@@
Evaluation Metrics.
We report image quality using SSIM, PSNR and LPIPS values, and throughput in training images per second. We take both forward and backward time into consideration of throughput. And note that throughput in images per second may differ from throughput in iterations per second, as one iteration includes the batch size number of images.
@@ 5.2 Performance and Memory Scaling
Figure 8: Even for the small “Train” scene, we achieve speedup with distributed training and larger batch sizes without compromising test PSNR. All configs are trained with 30K total images.
@@
Computation.
We evaluated how additional GPUs impact Grendel’s performance using both large-scale (rubble and train) and small-scale (Mip-Nerf360) datasets.
We used the Rubble scene to evaluate the training throughput. For this experiment we used 35 million Gaussians which have been trained to convergence. Because of the time required to render 4K images for this scene, we measured throughput for training over another 10,000 images times, and in Figure 8 we report throughput (in images per second) as we vary the number of GPUs (x-axis) and batch size (y-axis). We observe that we cannot render this scene with a single GPU (regardless of batch size) because of its memory requirements. Furthermore, both increasing the number of GPUs and increasing batch size yield performance improvements: performance increases from 5.55 images per second (4 GPUs, batch size 1) to 38.03 images per-second (32 GPUs, batch size 64) in our experiments. We provide additional details about this experiment in Appendix C.2.
Next, we use the $980\times 545$ resolution Train scene to evaluate image quality during scaling (this image is small enough to allow us to render the image fully). Our results in Figure 8 show that additional GPUs improve throughput while maintaining image quality when trained with the same total number of images. Notably, our 16-GPU setup with a batch size of 32 completes training on 30K images in just 2 minutes and 42.97 seconds, representing the state-of-the-art training speed to the best of our knowledge.
We observed in Figure 10 the same improvement in performance without degradation for other scenes taken from the Mip-Nerf360 dataset (first half) and the Tanks & Temple and Deep Blending datasets (second half), and using default hyperparameters from the 3DGS repository [8]. We train both configurations on the same number of images: 50k for Mip-NeRF 360 and 30k for the slightly smaller TT & DB datasets, to ensure convergence and a fair comparison. Additional details on this experiment are in Appendix C.2.
Memory Scaling.
Scaling the number of GPUs also increases available memory, and thus the amount of Gaussians that can be used to represent a scene. We evaluated this effect by using densification to add additional Gaussians until we ran out of memory. In Figure 12 we show number of Gaussians (in millions) when using batch size 1, 4 and 16, while varying the number of GPUs, and observe linear scaling. In §5.2.1 we show the utility of using additional Gaussians, and we provide additional details about this evaluation in Appendix C.2.
@@
Scaling to multiple GPUs allows Grendel to use a larger number of Gaussians to represent scenes. A larger number of Gaussians can capture fine grained scene details, and should thus be able to better reconstruct large-scale, high-resolution scenes. We evaluated this effect using three scenes: Rubble, MatrixCity Block_All, and Bicycle, and varied the number of Gaussians used by changing densification settings: we lowered the gradient norm threshold to initiate densification and reduced the threshold for splitting Gaussians instead of cloning until the densification mechanism produced the target number of Gaussians without manual interference.
We rendered Rubble and Matrix City Block_All using 16 GPUs and a batch size of 16, while we used 4 GPUs and a batch size of 4 for Bicycle. The difference in number of GPUs and batch sizes is due to differences in scene sizes: bicycle is much smaller than the other two datasets. In Figure 13 we show that image quality metrics (PSNR, SSIM and LPIPS) improve as we add more Gaussians. The red line in the Rubble and Matrix City Block_All graphs shows the number of Gaussians that can fit on a single GPU (Bicycle, being smaller, can be rendered on a single GPU). Figure 14 shows the rendered images as we scale Gaussians quantity, and demonstrates the quality improvements are human visible. These results demonstrate benefits of using more Gaussians, and demonstrate the necessity of multi-GPU 3DGS training systems like Grendel.
5.3 Ablation Study
Figure 10 and 12 illustrate that our load balancing techniques and increased batch size significantly improve training throughput on both 1080p Mip-NeRF360 dataset and 4K Rubble datasets, compared to the one GPU baseline and our straightforward distributed system with a conventional batch size of one and no load balancing. Although good speed can be achieved without load balancing, load balancing allows us to consistently achieve even higher throughput across various types and scales of scenes.
@@ 6 Related Works
Large-scale scene reconstruction. Prior work has proposed systems that use the divide-and-conquer approach to scale 3DGS to work with large scenes. VastGaussian [17], CityGaussisan [18], and Hierarchical Gaussian [9] divide large scenes into small regions, train on these smaller regions on a single GPU, and then merge the resulting images. Hierarchical Gaussian [9], Octree-GS [27] and Scaffold-GS [19] describe level-of-detail based approaches to adaptively reduce the number of Gaussians considered for particular scenes. None of these systems can consider a full large-scale scene directly and achieve the same quality as ours. Similar approaches [29, 33, 12] have also been used to scale NeRF, but they cannot directly apply to 3DGS which exhibits a different computation pattern as we discussed in Section 1.
Distributed training for neural networks. Existing work have exploited various types of parallelism to train neural networks across GPUs. These include data parallelism [14], tensor parallelism [28, 23], pipeline parallelism [7, 22] and FSDP [34, 26]. Several systems also support multiple types parallelism and/or aim to automatically partition the workload to optimize training speed [35, 31, 30]. However, as we discussed earlier (§1), neural network training and 3DGS have very different computation patterns. The former performs repeated layer-wise computation dominated by dense matrix multiply operations while the latter’s 3 stages training process is irregular and sparse. As a result, although Grendel’s distribution strategy may resemble those seen in existing work (e.g., FSDP [34]), the details are quite different.
Large Batch Size Training. Large batch training has bee widely adopted to improve the ML training performance and efficiency, but it has also been recognized by prior work [10] that increasing batch size can adversely impact model performance. This has led to the development of empirical rules, including linear scaling and learning rate warmum [4], and layer-wised adaptive rate scaling [32, 3]. Theoretical work [16, 20, 2] have also predicted and validated scaling laws for increasing batch sizes. Our work is inspired by these, but focuses on batch size scaling for 3DGS training.
7 Conclusion
We have described Grendel, a distributed 3D Gaussian training system. Our evaluation shows that Grendel can render high-resolution high-quality images from large scenes that are beyond the capabilities of existing approaches, and that it shows nearly perfect performance and memory scalability. We plan to open source Grendel so that the community can more easily experiment with 3DGS based reconstruction techniques for large scenes.
8 Acknowledgements
@@ A.1 Densification Process
Densification is the process by which 3DGS adds more Gaussians to improve details in a particular region. A Gaussian that shows significant position variance across training steps, might either be clones or split. The decision on whether to clone or split depends on whether their scale exceeds a threshold. Hyperparameter determine the start and stop iteration for densification, its frequency, the gradient threshold for initiating densification, and the scale threshold that determines whether to split or clone. To create more Gaussians, we need to increase the stop iteration and frequency, and decrease the gradient threshold for densification. If we aim to capture more details using smaller Gaussians, we should lower the scale threshold to split more Gaussians. The training process also includes pruning strategies such as eliminating Gaussians with low opacity and using opacity reset techniques to remove redundant Gaussians.
@@ A.3 Mixed Parallelism
Regarding the memory aspect, each Gaussian has independent transformed states, gradients, optimizer states, and parameters for each camera view. Therefore, we save these states together on the corresponding GPU that contains their parameters. And activation states like significant Z-buffers, auxiliary buffers for sorting and other functions, loss intermediate activations are managed pixel-wise along with the image distribution.
@@ A.4 Dynamic Unbalanced Workloads
Physical scenes are naturally sparse on a global scale. Different area has different densities of 3D gaussians (i.e sky and a tree). Thus, the intensity of rendering not only varies from pixel to pixel within an image but also differs between various images, leading to workloads unbalance.
@@
The different computational steps have distinct characteristics in terms of workload dynamicity. Even though, the rendering computation is dynamic and unbalanced; computation intensity for loss calculation remains consistent across pixels, and the view-dependent transformation maintains a uniform computational intensity across gaussians. Actually, render forward and backward have different pattern of unbalance and dynamicity. The computational complexity for the forward process scales with the number of 3DGS intersecting the ray. In contrast, the complexity of the backward process depends on Gaussians that contributed to color and loss before reaching opacity saturation, typically those on the first surface. Then, running time for render forward and backward, loss forward and backward have different dominating influence factors, and every step takes a significant amount of time.

### v2

@@ 4.1 Empirical Evidence of Independent Gradients
To see if the Independent Gradients Hypothesis holds in practice, we analyze the average per-parameter variance of the gradients in real-world settings. We plot the sparsity and variance of the gradients of the diffuse color parameters starting at pre-trained checkpoints on the “Rubble” dataset (Turki et al., 2022) against the batch size in Figure 6. We find that the inverse of the variance increases roughly linearly, then transitions into a plateau. We find this behavior in all three checkpoint iterations, representing early, middle, and late training stages. The initial linear increase of the precision suggests that gradients are roughly uncorrelated at batch sizes used in this work (up to 32) and supports the independent gradients hypothesis. While a single image may have sparse gradients; in a large batch, gradients overlap and become less sparse. They also grow more correlated, as camera with similar poses are expected to offer similar gradients.
Figure 6: Gradients are roughly uncorrelated in practice. On the “Rubble” dataset (Turki et al., 2022), the inverse of the average parameter variance increases linearly, then rises to a plateau, suggesting that the gradients are roughly uncorrelated initially but become less so as the batch size becomes large. Averaged over 32 random trials.
@@ 4.2 Empirical Testing of Proposed Scaling Rules
To empirically validate the proposed learning rate and momentum scaling rules, we train the “Rubble” scene up to iteration 15,000 using a batch size of 1. Then, we reset the Adam optimizer states and continue training with different batch sizes. We compare how well different learning rate and momentum scaling rules maintain a similar training trajectory when switching to larger batch sizes in Figure 12(Appendix B.3). Without loss of generality, we focus on the diffuse color parameters for this analysis. Figure 12(a) compares three different learning rate scaling rules $\in$ [constant, sqrt, linear] where only our proposed “sqrt” holds a high update cosine similarity and a similar update magnitude across different training batch sizes. Similarly, 12(b) shows our proposed exponential momentum scaling rule keeps update cosine similarity higher than the alternative which leaves the momentum coefficients unchanged.
@@ 5 Evaluation
Our evaluation aims to demonstrate Grendel’s scalability, showing both that it can render high-resolution images from large scenes, and that its performance scales with additional hardware resources. We also compare Grendel’s system-level parallelization with CityGaussian’s (Liu et al., 2024) divide-and-conquer approach in 5.3. The ablation study on dynamic load balancing and learning rate scaling strategies is presented in Appendix C.1.
@@ 5.1 Setting and Datasets
Experimental Setup. We conducted our evaluation in the Perlmutter GPU cluster NERSC . Each node we used was equipped with 4 A100 GPUs with 40GB of GPU memory, and interconnected with each other using 25GB/s NVLink per direction. Servers were connected to each other using a 200Gbps Slingshot network.
Datasets. We evaluate Grendel using the datasets and corresponding resolution settings shown in Table 1. Of these, Rubble and MatrixCity Block_All represent the large scale datasets that are out of reach for most existing 3DGS systems, while other datasets are commonly used in 3DGS papers. These datasets vary in area size and resolution to comprehensively test our system.
@@
Tanks & Temple (Knapitsch et al., 2017) 2 $\sim$ 1K 251 to 301 1/8 of all images
DeepBlending (Hedman et al., 2018) 2 $\sim$ 1K 225 to 263 1/8 of all images
Mip-NeRF 360 (Barron et al., 2022) 9 1080P 100 to 330 1/8 of all images
Rubble (Turki et al., 2022) 1 $4591\times 3436$ 1657 official test set
MatrixCity Block_All (Li et al., 2023) 1 1080P 5620 official test set
@@
Evaluation Metrics. We report image quality using SSIM, PSNR and LPIPS values, and throughput in training images per second. We take both forward and backward time into consideration of throughput. And note that throughput in images per second may differ from throughput in iterations per second, as one iteration includes the batch size number of images.
@@ 5.2 Performance and Memory Scaling
Figure 8: Even for the small ’Train’ scene, our distributed training with larger batch sizes achieves speedup without compromising test PSNR. All configs train for 30K total images.
@@
Computation. We evaluated how additional GPUs impact Grendel’s performance using both large-scale (Rubble) and small-scale (Train and Mip-Nerf360) datasets.
We used the Rubble scene to evaluate the training throughput. For this experiment we used 35 million Gaussians which have been trained to convergence. Because of the time required to render 4K images for this scene, we measured throughput for training over another 10,000 images times, and in Figure 8 we report throughput (in images per second) as we vary the number of GPUs (x-axis) and batch size (y-axis). We observe that we cannot render this scene with a single GPU (regardless of batch size) because of its memory requirements. Furthermore, both increasing the number of GPUs and increasing batch size yield performance improvements: performance increases from 5.55 images per second (4 GPUs, batch size 1) to 38.03 images per-second (32 GPUs, batch size 64).
Next, we use the $980\times 545$ resolution Train scene to evaluate both throughput and image quality during scaling. Due to its small, low-resolution nature, it can be trained from scratch for each experiment. Our results in Figure 8 show that additional GPUs improve throughput while maintaining image quality when trained with the same total number of images. Notably, our 16-GPU setup with a batch size of 32 completes training on 30K images in just 2 minutes and 42.97 seconds, representing the state-of-the-art training speed to the best of our knowledge.
As shown in Figure 10, we also achieve a 3x to 4x speed up using 4 GPU and a batch size of 4, without PSNR degradation across 13 scenes from the Mip-Nerf360 dataset (first half) and the Tanks & Temple and Deep Blending datasets (second half). We use default hyperparameters from the 3DGS repository (Kerbl et al., 2023). We train on the same number of images: 50k for Mip-NeRF 360 and 30k for the slightly smaller TT & DB datasets, to ensure convergence and a fair comparison.
Memory Scaling. Scaling the number of GPUs increases memory, allowing more Gaussians to represent a scene. We tested this by adding Gaussians through densification until we ran out of memory. Figure 15 (In Appendix C.3) illustrates the number of Gaussians that Grendel can accommodate (in millions) with batch sizes of 1, 4, and 16, as the number of GPUs increases. The results demonstrate linear scaling. In §5.2.1 we show the utility of using additional Gaussians.
We provide additional details about experiments from this section in Appendix C.3.
@@
Scaling to multiple GPUs allows Grendel to use a larger number of Gaussians to represent scenes. A larger number of Gaussians can capture fine grained scene details, and should thus be able to better reconstruct large-scale, high-resolution scenes. We evaluated this effect using three scenes: Rubble, MatrixCity Block_All, and Bicycle, and varied the number of Gaussians used by changing densification settings: we lowered the gradient norm threshold to initiate densification and reduced the threshold for splitting Gaussians instead of cloning until the densification mechanism produced the target number of Gaussians without manual interference. We rendered Rubble and Matrix City Block_All using 16 GPUs and a batch size of 16, while we used 4 GPUs and a batch size of 4 for Bicycle. The difference in number of GPUs and batch sizes is due to differences in scene sizes: bicycle is much smaller than the other two datasets.
In Figure 11, we show that image quality metrics (PSNR, SSIM and LPIPS) improve as we add more Gaussians. The red line in the Rubble and Matrix City Block_All graphs shows the number of Gaussians that can fit on a single GPU (Bicycle, being smaller, can be rendered on a single GPU). Figure 13 shows the rendered images as we scale Gaussians quantity, and demonstrates the quality improvements are human visible. These results demonstrate benefits of using more Gaussians, and demonstrate the necessity of multi-GPU 3DGS training systems like Grendel.
5.3 Comparison to CityGaussian
We evaluate Grendel’s system-level parallelization against CityGaussian(Liu et al., 2024), the open-source 3DGS divide-and-conquer solution. Since CityGaussian is not designed for high-resolution reconstruction, we perform these experiments on downsampled scenes: Rubble (4x downsampled) (Turki et al., 2022), Building (4x downsampled)(Turki et al., 2022), and MatrixCity Block_All (downsampled to a width of 1600 pixels)(Li et al., 2023). All experiments are conducted on 4 A100s. We provide two comparison points for CityGaussian. CityGS Official uses the authors’ official script  [^1] https://github.com/DekuLiuTesla/CityGaussian/tree/main/scripts, yielding PSNR numbers comparable to those reported in (Liu et al., 2024). CityGS Official experiments train Rubble, Building, and MatrixCity Block_All on a total of 300000, 630000, and 1110000 images, respectively. CityGS 200K-images experiments use the same configuration as CityGS Official, except for training each scene with a total of only 200000 images, in order to compare the convergence rate with Grendel. In Grendel 200K-images, each scene is trained using Grendel for a total of 200000 images, matching the count in CityGS 200K-images.
Method Rubble Building MatrixCity Block_All
PSNR SSIM LPIPS Total Time PSNR SSIM LPIPS Total Time PSNR SSIM LPIPS Total Time
CityGS Official 25.88 0.813 0.231 2.88 hrs 22.14 0.784 0.241 4.57 hrs 27.41 0.864 0.205 8.25 hrs
CityGS 200K-images 25.40 0.796 0.249 2.18 hrs 20.32 0.725 0.299 2.22 hrs 23.68 0.701 0.422 3.60 hrs
Grendel 200K-images (ours) 27.39 0.859 0.195 0.85 hrs 22.69 0.778 0.242 0.90 hrs 27.33 0.859 0.205 1.22 hrs
[…]
As shown in Table 2, Grendel achieves test PSNR comparable to, or surpassing those of CityGS Official. Grendel is much faster—achieving 3x-6.7x speed improvements over CityGS Official. Compared to CityGS 200K-images, Grendel demonstrates greater training efficiency in both convergence rate and training time. Our experience also shows that Grendel is simpler to use. CityGS requires running several separate procedures each of which requires hyperparameter tuning. By contrast, running Grendel over multiple GPUs requires similar efforts as the original 3DGS.
@@ 6 Related Works
Large-scale scene reconstruction. Prior works have proposed the divide-and-conquer approach to scale 3DGS to work with large scenes. VastGaussian (Lin et al., 2024), CityGaussian (Liu et al., 2024), and Hierarchical Gaussian (Kerbl et al., 2024) divide large scenes into small regions and train each region separately, and then merge the resulting sub-models. Hierarchical Gaussian (Kerbl et al., 2024), Octree-GS (Ren et al., 2024) and CityGaussisan (Liu et al., 2024) describe level-of-detail based approaches to adaptively reduce the number of Gaussians considered for distant objects. Grendel’s system-level parallelization is complementary to these algorithmic innovations. For example, the initial coarse training step in both CityGaussisan (Liu et al., 2024), and Hierarchical Gaussian (Kerbl et al., 2024) can utilize Grendel for multi-GPU acceleration. Similar methods (Turki et al., 2022; Yuanbo et al., 2022; Li et al., 2024b) have been employed to scale NeRF, but they are not directly applicable to 3DGS due to its distinct computation pattern, as discussed in Section 1.
Distributed Training for 3DGS. DOGS (Chen & Lee, 2024) modifies training with ADMM distributed optimization, averaging Gaussians shared across partition boundaries every 100 iterations. This method enables asynchronous training, which can potentially impact convergence rate. In contrast, Grendel preserves the original 3DGS algorithm, maintaining the same convergence characteristics as the single-GPU version. RetinaGS (Li et al., 2024a) also retains the original 3DGS algorithm but employs a distinct parallelization strategy. In RetinaGS, each GPU renders the entire image using its local partition of Gaussians, with the rendered outputs subsequently merged. However, this approach results in redundant computations, as many GPUs render pixels beyond the opacity saturation depth unnecessarily. Distributed training for neural networks is further discussed in Appendix D.
Large Batch Size Training. Large batch training has been widely adopted to improve the ML training performance and efficiency, but it has also been recognized by prior work (Keskar et al., 2017) that increasing batch size can adversely impact model performance. This has led to the development of empirical rules for neural networks training, including linear scaling and learning rate warmum for SGD (Goyal et al., 2017), square root scaling for Adam (Malladi et al., 2022) and layer-wised adaptive rate scaling (You et al., 2020; Ginsburg et al., 2018). The scaling rules in our work are inspired by these, but focuses on batch size scaling for 3DGS.
Reproducibility Statement
We provide hyperparameter details in Appendix C.3. We will also release code for training, rendering, and evaluation. We will release pre-trained model checkpoints for Rubble and MatrixCity datasets to reproduce the results reported in the paper.
Acknowledgements
@@ A.1 Densification Process
Densification is the process by which 3DGS adds more Gaussians to improve details in a particular region. A Gaussian that shows significant position variance across training steps, might either be clones or split. The decision on whether to clone or split depends on whether their scale exceeds a threshold. Hyperparameters determine the start and stop iteration for densification, its frequency, the gradient threshold for initiating densification, and the scale threshold that determines whether to split or clone. To create more Gaussians, we need to increase the stop iteration and frequency, and decrease the gradient threshold for densification. If we aim to capture more details using smaller Gaussians, we should lower the scale threshold to split more Gaussians. The training process also includes pruning strategies such as eliminating Gaussians with low opacity and using opacity reset techniques to remove redundant Gaussians.
@@ A.3 Mixed Parallelism
Regarding the memory aspect, each Gaussian has independent transformed states, gradients, optimizer states, and parameters. Therefore, we save these states together on the corresponding GPU that contains their parameters. And activation states like significant Z-buffers, auxiliary buffers for sorting and other functions, loss intermediate activations are managed pixel-wise along with the image distribution.
@@ A.4 Dynamic Unbalanced Workloads
Physical scenes are naturally sparse on a global scale. Different areas have different densities of 3D gaussians (i.e sky and a tree). Thus, the intensity of rendering not only varies from pixel to pixel within an image but also differs between various images, leading to workloads unbalance. Figure 5 shows the differences in render intensity across the image.
@@
The different computational steps have distinct characteristics in terms of workload dynamicity. Even though, the rendering computation is dynamic and unbalanced; computation intensity for loss calculation remains consistent across pixels, and the view-dependent transformation maintains a uniform computational intensity across gaussians. Actually, render forward and backward have different patterns of unbalance and dynamicity. The computational complexity for the forward process scales with the number of 3DGS intersecting the ray. In contrast, the complexity of the backward process depends on Gaussians that contributed to color and loss before reaching opacity saturation, typically those on the first surface. Then, running time for render forward and backward, loss forward and backward have different dominating influence factors, and every step takes a significant amount of time.